OpenStudy (cicilybailey14):

Fan+Medal! Do the data in the table represent a direct variation or an inverse variation? Write an equation to model the data in the table. x y 3 24 9 8 12 6 direct; y = 8x inverse; xy = 8 inverse; xy = 72 direct; y = 72x

OpenStudy (michele_laino):

I think data that represent an inverse variation, have to check this equation: x*y=k. So substituting your numerical data, we have: x*y=24*3=9*8=12*6=72

OpenStudy (michele_laino):

so k=72

OpenStudy (cicilybailey14):

I wasnt sure ....I have this slide that might help?

OpenStudy (michele_laino):

yes it can!

OpenStudy (cicilybailey14):

So is it c?

OpenStudy (michele_laino):

that's right!

OpenStudy (cicilybailey14):

Alright thanks! :D

OpenStudy (michele_laino):

thanks!:)